# Generative AI for Low-Carbon Artificial Intelligence of Things

### v1

@@ Abstract
By integrating Artificial Intelligence (AI) with the Internet of Things (IoT), Artificial Intelligence of Things (AIoT) has revolutionized many fields. However, AIoT is facing the challenges of energy consumption and carbon emissions due to the continuous advancement of mobile technology. Fortunately, Generative AI (GAI) holds immense potential to reduce carbon emissions of AIoT due to its excellent reasoning and generation capabilities. In this article, we explore the potential of GAI for carbon emissions reduction and propose a novel GAI-enabled solution for low-carbon AIoT. Specifically, we first study the main impacts that cause carbon emissions in AIoT, and then introduce GAI techniques and their relations to carbon emissions. We then explore the application prospects of GAI in low-carbon AIoT, focusing on how GAI can reduce carbon emissions of network components. Subsequently, we propose a Large Language Model (LLM)-enabled carbon emission optimization framework, in which we design pluggable LLM and Retrieval Augmented Generation (RAG) modules to generate more accurate and reliable optimization problems. Furthermore, we utilize Generative Diffusion Models (GDMs) to identify optimal strategies for carbon emission reduction. Simulation results demonstrate the effectiveness of the proposed framework. Finally, we insightfully provide open research directions for low-carbon AIoT.
@@ I Introduction
Currently, Artificial Intelligence of Things (AIoT) is ushering in a new era of digital economy by supporting the technological revolution in many fields[1], such as smart healthcare and smart agriculture[2]. However, the impact of AIoT on energy consumption and carbon emissions is a topic of concern[3]. Specifically, the advent of transformative technologies such as AI-Generated Content (AIGC), the Internet of Things (IoT), and Metaverses has led to a significant surge in data volume within AIoT[4]. According to research conducted by Transforma Insights, the global deployment of edge devices is projected to rise from 2.7 billion to 7.8 billion in the next decade. Furthermore, the broader category of IoT-connected devices is expected to surpass 30 billion worldwide by 2025, and the mobile data traffic of a mobile device will reach 257.1 GB per month by 2030, which is a substantial increase of 50 times compared to the data volume in 2010[4]. Hence, the rapid growth in power consumption of edge loads and the scarcity of energy resources pose significant energy challenges to AIoT, resulting in new environmental impacts. It is worth noting that low energy consumption and low carbon emissions are related but distinct concepts in terms of environmental impacts. Specifically,
@@
Inspired by the revolutionary capabilities of GAI, GAI-driven solutions hold the immense potential to optimize energy consumption and reduce carbon emissions in AIoT. In addition, GAI can cope with the dynamic changes of network conditions and adaptively adjust optimal strategies without retraining, avoiding additional carbon footprints. Therefore, GAI opens up new avenues to achieve low-carbon AIoT. Unlike traditional green mobile networks, which primarily focus on reducing energy consumption and enhancing energy efficiency[5], low-carbon AIoT enabled by GAI focuses on utilizing GAI to minimize carbon emissions and promotes sustainable practices across the entire network ecosystem, which has the following potential characteristics:
@@
Fig. 1 presents recent advances in integrating GAI with intelligent networks and related concepts of carbon emission goals, including low-carbon, carbon-free, carbon-neutral, and net-zero carbon. To the best of our knowledge, this is the first work that systemically provides forward-looking research on the potential of GAI enabling low-carbon AIoT, which is the first step toward carbon-neutral and net-zero carbon paradigms. Our main contributions are summarized as follows:
  • We first investigate the main carbon emission impacts of mobile networks, and then briefly discuss the limitations of DAI in carbon emission reduction, and systematically introduce GAI techniques, including their features and abilities to reduce carbon emissions.
@@
  • We propose an LLM combined with Retrieval Augmented Generation (RAG) framework for carbon emission optimization, in which we adopt Generative Diffusion Models (GDMs) to identify optimal strategies. Simulation results of a real carbon emission optimization case study demonstrate the effectiveness of the proposed framework.
@@ II-A Carbon Emission from Mobile Networks
C2. Computational impact on carbon emissions: While benefiting human productivity and efficiency, the large-scale use of computational devices has led to the explosion of data and computation in mobile networks, resulting in huge carbon emissions. For example, there were 7.7 billion mobile phones in use worldwide in 2020, producing about 580 million tons of carbon emissions, equivalent to about 1% of total global emissions. In mobile networks, the limited power capacity of edge devices may also present challenges in affording substantial computation energy consumption required for computational-intensive tasks, especially for AI model training and inference[4]. For instance, the energy consumption for training a ResNet-110 model [^4] https://builtin.com/artificial-intelligence/resnet-architecture on the NVIDIA Jetson TX2 platform amounts to approximately 8 × 105 Joules of energy[5].
@@ II-B Discriminative AI in Carbon Emission Reduction
As a class of AI that aims to distinguish between different classes in a given dataset, DAI has been utilized in many promising sectors to reduce carbon emissions, such as renewable energy harvest[3], carbon capture[9], and energy management[3]. For example, the authors in [3] proposed a machine learning model called Deep Energy to effectively coordinate the working state of 5G cells and avoid carbon efficiency traps. However, due to the dynamic and heterogeneous nature of AIoT[2], DAI has obvious limitations in terms of carbon reduction in AIoT.
  • Limited applicability: The efficacy of DAI in carbon emission reduction heavily depends on the specific applications[9], such as renewable energy integration, transportation optimization, and smart grid management. When new applications emerge, the existing DAI models need to be retrained, resulting in huge energy consumption and carbon emissions.
@@ II-C Generative AI Technologies and their Relations to Carbon Emissions
  • Generative Adversarial Networks (GANs): GANs consist of generator and discriminator networks, where the generator network aims to generate new data and the discriminator network aims to distinguish synthetic data from real data[2]. Since the two networks engage in iterative training and competition, GANs possess data generation and discrimination capabilities. In energy optimization, GANs, such as BiLSTM-CNN-GAN [^5] https://typeset.io/questions/what-are-the-bilstm-cnn-gan-algorithm-3wt75zl0t7, can predict energy consumption and carbon emissions, facilitating efficient resource management and planning.
  • Retrieval Augmented Generation: RAG is an advanced technique for enhancing the reliability and accuracy of GAI models by retrieving facts from an external knowledge base[10]. Furthermore, RAG can augment user prompts by adding relevant retrieved data in context to allow LLMs to generate accurate answers. In energy optimization, LLMs enabled by RAG can generate accurate energy production strategies by accessing external databases, such as documents about energy production.
  • Generative Diffusion Models: GDMs consist of forward diffusion and denoising processes inspired by non-equilibrium thermodynamics theory[2], gradually transforming initial random samples into the target output through several iterative denoising steps. With the incredible capability of image generation, GDMs have the potential to be applied to optimize energy-efficient image generation tasks, ensuring a more sustainable use of computing resources.
  • Other GAI techniques: Variational Autoencoders (VAEs) can represent data in a probabilistic latent space, which enhances accuracy in short-term energy forecasting and optimizes energy usage for efficiency. Flow-based Generative Models (FGMs) facilitate data generation by transforming input data distributions from simple to complex through a series of reversible transformations, which can be potentially applied to simulate the energy consumption of AIoT.
@@ III-A1 GAI-driven Variable Renewable Energy (VRE) harvesting
In EI, the integration of renewable energy within residential areas can enhance energy supplies for community-level mobile edge networks, which reduces carbon emissions and mitigates air pollution. In [6], the authors highlighted the significant role of GAI in advancing renewable and sustainable energy technologies. Specifically, by analyzing real-time data on environmental conditions and network demands, GAI can intelligently adjust energy harvesting mechanisms[6]. For instance, GAI can optimize the positions of solar panels installed for base stations to capture the maximum amount of solar energy during sunny days with ample sunlight. In addition, GAI can efficiently estimate renewable and sustainable electricity production using spatial and temporal data from other renewable energy sources, such as biomass and onshore wind energy. This estimation can be used to schedule network and computing workloads to reduce reliance on carbon-based power generation. Therefore, GAI-driven VRE harvesting holds immense potential to mitigate environmental impacts and enhance energy accessibility in mobile networks.
@@ III-A2 GAI-driven energy routing for Vehicle-to-Grid (V2G)
As a typical EI scenario, V2G technology integrates Electric Vehicles (EVs) into the power grid. However, the main power source of EVs does not rely only on the power grid, but also on a variety of other energy sources, such as renewable energy sources[11]. In this case, the energy is transmitted throughout the networked EI scenario, like information routing on the Internet[11]. It is worth noting that power losses and carbon reduction are mutually exclusive in energy transmission[11]. As a result, how to design a proper energy routing strategy to simultaneously satisfy these two targets is significant. GAI has been applied in designing routing strategies[12]. Combined with DRL, GAI can analyze complex data sets to optimize routing decisions.
@@ III-B1 GAI-driven ICT and cooling system management
The electricity from ICT and cooling systems accounts for about 86% of the total energy consumption of data center networks[13]. In the data center network, the ICT system generates heat during operation, and the cooling system is designed to dissipate this heat to maintain a suitable for the equipment. Thus, the effective management of both ICT and cooling systems is essential to enhance energy efficiency and reduce carbon emissions in data center networks[13]. GAI is expected to optimize the management of ICT and cooling systems in data center networks. For instance, by analyzing real-time data from ICT and cooling systems, GAI can predict their equipment failures in advance[2], which optimizes the corresponding maintenance activities and reduces energy wastage.
@@ III-B2 GAI-driven network optimization
Figure 3: The LLM-enabled carbon emission optimization framework enabled by RAG. In the proposed framework, RAG consists of three key components, i.e., database, retrieval, and decision-making, enabling accurate and reliable carbon emission optimization problems. Then, GDMs can be employed to generate optimal strategies, where the motivations and specific process of utilizing GDMs for determining optimal strategies are introduced in [14].
@@ III-C2 GAI-driven optimal IRS deployment
IV LLM-enabled Carbon Emission Optimization Framework Enabled by RAG
In this section, we propose an LLM-enabled carbon emission optimization framework enabled by RAG. We conduct a case study on carbon emission optimization for mobile AIGC task offloading in a metaverse environment and utilize GDMs to generate optimal strategies.
@@ IV-A Motivation
Carbon emission optimization is a significant approach for minimizing environmental impacts from AIoT. It refers to optimizing specially various sectors of mobile systems, such as data transfer energy efficiency[3], data centers[13], and task offloading[15]. Inspired by exceptional decision-making capability of LLMs, we propose an LLM-enabled carbon emission optimization framework enabled by RAG for AIoT. By interpreting the network environment, the proposed framework can automatically formulate significant carbon emission optimization problems through simple interactions with network designers. With RAG, the framework can significantly lower the risk of human errors, improve the accuracy of problem formulation, and speed up the design process by fusing the carbon emission reduction knowledge learned from the comprehensive knowledge base[10].
@@ IV-B Framework Design
Upon an optimization problem generated by the LLM, the network designer can simply summarize and correct the formulated optimization problem. Since GDMs show superior performance in handling high-dimensional and complex optimization problems[14], the network designer can utilize GDMs to generate optimal strategies and implement these strategies in real-world scenarios, thereby effectively achieving carbon emission reduction in AIoT. The technical principle and specific process of GDMs for solving optimization problems can be found in [14].
@@ IV-C Case Study: Mobile AIGC Task Offloading in a Metaverse Environment
IV-C1 Framework configuration
In our experiments, we call the GPT 4 model through the OpenAI API to implement the pluggable LLM module, and the RAG module is built on top of LangChain [^6] https://www.langchain.com/.
IV-C2 Scenario description
In the scenario of mobile AIGC services in a metaverse environment, users request AIGC services from edge servers, such as personalized avatars that provide users with immersive experiences in the metaverse. Edge servers, powered by renewable energy sources[15], fine-tune pre-trained AIGC models and execute inferences to enhance the quality of immersive experiences for the users. To reduce service latency, AIGC tasks can be collaboratively executed by the edge servers and users. In particular, we consider a user and an edge server in this scenario. By optimizing bandwidth and transmit power of the edge server, the goal is to minimize the carbon emissions of an AIGC task through the offloading mechanism while ensuring high-quality AIGC services in the metaverse setting.

### v2

@@ Abstract
By integrating Artificial Intelligence (AI) with the Internet of Things (IoT), Artificial Intelligence of Things (AIoT) has revolutionized many fields. However, AIoT is facing the challenges of energy consumption and carbon emissions due to the continuous advancement of mobile technology. Fortunately, Generative AI (GAI) holds immense potential to reduce carbon emissions of AIoT due to its excellent reasoning and generation capabilities. In this article, we explore the potential of GAI for carbon emissions reduction and propose a novel GAI-enabled solution for low-carbon AIoT. Specifically, we first study the main impacts that cause carbon emissions in AIoT, and then introduce GAI techniques and their relations to carbon emissions. We then explore the application prospects of GAI in low-carbon AIoT, focusing on how GAI can reduce carbon emissions of network components. Subsequently, we propose a Large Language Model (LLM)-enabled carbon emission optimization framework, in which we design pluggable LLM and Retrieval Augmented Generation (RAG) modules to generate more accurate and reliable optimization problems. Furthermore, we utilize Generative Diffusion Models (GDMs) to identify optimal strategies for carbon emission reduction. Numerical results demonstrate the effectiveness of the proposed framework. Finally, we insightfully provide open research directions for low-carbon AIoT.
@@ I Introduction
Currently, Artificial Intelligence of Things (AIoT) is ushering in a new era of the digital economy by supporting the technological revolution in many fields[1], such as smart healthcare and smart agriculture[2]. However, the impact of AIoT on energy consumption and carbon emissions is a topic of concern[3]. Specifically, the advent of transformative technologies such as AI-Generated Content (AIGC), the Internet of Things (IoT), and Metaverse has led to a significant surge in data volume within AIoT[4]. According to research conducted by Transforma Insights, the global deployment of edge devices is projected to rise from 2.7 billion to 7.8 billion in the next decade. Furthermore, the broader category of IoT-connected devices is expected to surpass 30 billion worldwide by 2025, and the mobile data traffic of a mobile device will reach 257.1 GB per month by 2030, which is a substantial increase of 50 times compared to the data volume in 2010[4]. Hence, the rapid growth in power consumption of edge loads and the scarcity of energy resources pose significant energy challenges to AIoT, resulting in new environmental impacts. It is worth noting that low energy consumption and low carbon emissions are related but distinct concepts in terms of environmental impacts. Specifically,
@@
Current research commonly focuses on using Discriminative AI (DAI) to reduce network carbon emissions[3, 7]. However, DAI which focuses primarily on analyzing or classifying existing data has a poor capability of adapting to the dynamic environment of AIoT. Inspired by the revolutionary capabilities of GAI, GAI-driven solutions hold the immense potential to optimize energy consumption and reduce carbon emissions in AIoT. In addition, GAI can cope with the dynamic changes of network conditions and adaptively adjust optimal strategies without retraining, avoiding additional carbon footprints. Therefore, GAI opens up new avenues to achieve low-carbon AIoT. Unlike traditional green mobile networks, which primarily focus on reducing energy consumption and enhancing energy efficiency, low-carbon AIoT enabled by GAI focuses on utilizing GAI to minimize carbon emissions and promotes sustainable practices across the entire network ecosystem, which has the following potential characteristics:
@@
Figure 1 presents recent advances in integrating GAI with intelligent networks and related concepts of carbon emission goals, including low-carbon, carbon-free, carbon-neutral, and net-zero carbon. To the best of our knowledge, this is the first work that systemically provides forward-looking research on the potential of GAI enabling low-carbon networks, which is the first step toward carbon-neutral and net-zero carbon paradigms. Our main contributions are summarized as follows:
  • We first investigate the main carbon emission impacts of mobile networks, then briefly discuss the limitations of DAI in carbon emission reduction, and systematically introduce GAI techniques, including their features and abilities to reduce carbon emissions.
@@
  • We propose an LLM framework combining Retrieval Augmented Generation (RAG) for carbon emission optimization, where we design pluggable LLM and RAG modules that rely on knowledge bases and context memory to generate carbon emission optimization problems.
  • We adopt Generative Diffusion Models (GDMs) to identify optimal strategies for carbon emissions. Simulation results of a real carbon emission optimization case study demonstrate the effectiveness of the proposed framework.
@@ II-A Carbon Emission from Mobile Networks
C2. Computational impact on carbon emissions: While benefiting human productivity and efficiency, the large-scale use of computational devices has led to the explosion of data and computation in mobile networks, resulting in huge carbon emissions. For example, there were 7.7 billion mobile phones in use worldwide in 2020, producing about 580 million tons of carbon emissions, equivalent to about 1% of total global emissions. In mobile networks, the limited power capacity of edge devices may also present challenges in affording substantial computation energy consumption required for computational-intensive tasks, especially for AI model training and inference[4]. For instance, the energy consumption for training a ResNet-110 model [^4] https://builtin.com/artificial-intelligence/resnet-architecture on the NVIDIA Jetson TX2 platform amounts to approximately 8 × 105 Joules of energy.
@@ II-B Discriminative AI in Carbon Emission Reduction
As a class of AI that aims to distinguish between different classes in a given dataset, DAI has been utilized in many specific tasks for reducing carbon emissions, such as renewable energy harvest[3], carbon capture[7], and energy management[3]. For example, the authors in [3] proposed a machine learning model to effectively coordinate the working state of 5G cells and avoid carbon efficiency traps. However, due to the dynamic and heterogeneous nature of AIoT[2], DAI has obvious limitations in terms of carbon reduction:
  • Limited applicability: The efficacy of DAI in carbon emission reduction heavily depends on specific applications[7], such as renewable energy integration, transportation optimization, and smart grid management. When new applications emerge, the existing DAI models need to be retrained, resulting in huge energy consumption and carbon emissions.
@@ II-C Generative AI Technologies and their Relations to Carbon Emissions
  • Generative Adversarial Networks (GANs): GANs consist of generator and discriminator networks, where the generator network aims to generate new data and the discriminator network aims to distinguish synthetic data from real data[2]. Since the two networks engage in iterative training and competition, GANs possess data generation and discrimination capabilities. For carbon emission reduction, GANs, such as BiLSTM-CNN-GAN [^5] https://typeset.io/questions/what-are-the-bilstm-cnn-gan-algorithm-3wt75zl0t7, can predict energy consumption and carbon emissions, facilitating efficient resource management and planning.
  • Retrieval Augmented Generation: RAG is an advanced technique for enhancing the reliability and accuracy of GAI models by retrieving facts from an external knowledge base[6], which can augment user prompts by adding relevant retrieved data in context to allow LLMs to generate accurate answers. Especially, LLMs supported by RAG can generate accurate carbon emission optimization strategies by accessing external databases, such as documents about carbon emission reduction.
  • Generative Diffusion Models: GDMs consist of forward diffusion and denoising processes inspired by non-equilibrium thermodynamics theory[2], gradually transforming initial random samples into the target output through several iterative denoising steps. With the incredible capability of image generation, GDMs have the potential to be applied to optimize image generation tasks, ensuring a more sustainable use of computing resources and reducing carbon emissions.
  • Other GAI techniques: Variational Autoencoders (VAEs) can represent data in a probabilistic latent space, which enhances accuracy in short-term energy forecasting and optimizes energy usage for carbon emission reduction. Flow-based Generative Models (FGMs) facilitate data generation by transforming input data distributions from simple to complex through a series of reversible transformations, which can be potentially applied to predict weather patterns, ensuring that as much renewable energy as possible is harvested to reduce carbon emissions.
@@ III-A1 GAI-driven Variable Renewable Energy (VRE) harvesting
In EI, the integration of renewable energy within residential areas can enhance energy supplies for community-level mobile edge networks, which reduces carbon emissions and mitigates air pollution. In [8], the authors highlighted the significant role of GAI in advancing renewable and sustainable energy technologies. Specifically, by analyzing real-time data on environmental conditions and network demands, GAI can intelligently adjust energy harvesting mechanisms[8]. For instance, GAI can optimize the positions of solar panels installed for base stations to capture the maximum amount of solar energy during sunny days with ample sunlight. In addition, GAI can efficiently estimate renewable and sustainable electricity production using spatial and temporal data from other renewable energy sources, such as biomass and onshore wind energy. This estimation can be used to schedule network and computing workloads to reduce reliance on carbon-based power generation. Thus, GAI-driven VRE harvesting holds immense potential to mitigate environmental impacts and enhance energy accessibility in mobile networks.
@@ III-A2 GAI-driven energy routing for Vehicle-to-Grid (V2G)
As a typical EI scenario, V2G technology integrates Electric Vehicles (EVs) into the power grid. However, the main power source of EVs does not rely only on the power grid, but also on a variety of other energy sources, such as renewable energy sources[11]. In this case, the energy is transmitted throughout the networked EI scenario, like information routing on the Internet[11]. It is worth noting that power losses and carbon reduction are mutually exclusive in energy transmission[11]. As a result, how to design a proper energy routing strategy to simultaneously satisfy these two targets is significant. GAI has been applied in designing routing strategies[12]. Combined with DRL, GAI can analyze complex data sets to optimize routing decisions for reducing unnecessary carbon emissions.
@@ III-B1 GAI-driven ICT and cooling system management
The electricity from ICT and cooling systems accounts for about 86% of the total energy consumption of data center networks[13]. In the data center network, the ICT system generates heat during operation, and the cooling system is designed to dissipate this heat to maintain a suitable for the equipment. Thus, the effective management of both ICT and cooling systems is essential to enhance energy efficiency and reduce carbon emissions in data center networks[13]. GAI is expected to optimize the management of ICT and cooling systems in data center networks. For instance, by analyzing real-time data from ICT and cooling systems, GAI can predict their equipment failures in advance[2], optimizing the corresponding maintenance activities and reducing energy wastage.
@@ III-B2 GAI-driven network optimization
Figure 3: The LLM-enabled carbon emission optimization framework supported by RAG. In the proposed framework, RAG consists of three key components, i.e., database, retrieval, and decision-making, enabling accurate and reliable carbon emission optimization problems. Then, GDMs can be employed to generate optimal strategies, where the motivations and the specific process of utilizing GDMs for determining optimal strategies are introduced in [5].
@@ III-C2 GAI-driven optimal IRS deployment
IV LLM-enabled Carbon Emission Optimization Framework Supported by RAG
In this section, we propose an LLM-enabled carbon emission optimization framework supported by RAG. We conduct a case study on carbon emission optimization for mobile AIGC task offloading in a metaverse environment and utilize GDMs to generate optimal strategies.
@@ IV-A Motivation
Carbon emission optimization is a significant approach for minimizing environmental impacts from AIoT. It specifically refers to optimizing various sectors of mobile systems, such as data transfer energy efficiency[3], data centers[13], and task offloading[14]. Inspired by the exceptional decision-making capability of LLMs, we propose an LLM-enabled carbon emission optimization framework supported by RAG. By interpreting the network environment, the proposed framework can automatically formulate significant carbon emission optimization problems through simple interactions with network designers. With the support of RAG[6], the framework can significantly lower the risk of human errors, improve the accuracy of problem formulation[15], and speed up the design process by fusing the carbon emission reduction knowledge learned from the comprehensive knowledge base. Compared with the existing energy management strategies[14], the generated strategy from the problem formulated by the RAG-supported LLM agent is more comprehensive and practical.
@@ IV-B Framework Design
Upon the optimization problem generated by the LLM, the network designer can simply collate the generated optimization problem according to the requirements. Specifically, the network designer can customize subjective constraints and determine experimental parameters based on real scenarios. This process is almost burdenless. Since GDMs show superior performance in handling high-dimensional and complex optimization problems[5], the network designer can leverage GDMs to generate optimal strategies and implement these strategies in real-world scenarios, thereby effectively achieving carbon emission reduction in AIoT. The technical principle and specific process of GDMs for solving optimization problems can be found in [5]. Note that the trained GDM can adapt to different states of AIoT systems to generate the optimal strategy for carbon emission reduction.
In summary, the proposed LLM-enabled carbon emission optimization framework can help the network designer consider more factors for carbon emission reduction. In addition, RAG assists LLM agents to perform accurate inference and reduce the carbon emissions caused by inference tasks of LLMs, and prompt engineering can be applied within RAG to further enhance the interaction between the network designer and the LLM agent, allowing for precise information retrieval and generation based on finely tuned prompts.
@@ IV-C Case Study: Mobile AIGC Task Offloading in a Metaverse Environment
IV-C1 Scenario description
In the scenario of mobile AIGC services in a metaverse environment, users request AIGC services from edge servers, such as personalized avatars that provide users with immersive experiences in the metaverse. Edge servers, powered by renewable energy sources[14], fine-tune pre-trained AIGC models and execute inferences to enhance the quality of immersive experiences for the users. To reduce service latency, AIGC tasks can be collaboratively executed by the edge servers and users, where the intermediate results of AIGC tasks are sent to users by the edge servers. In particular, we consider a user and an edge server in this scenario. By optimizing the bandwidth and transmit power of the edge server, the goal is to minimize the carbon emissions of an AIGC task through the offloading mechanism while ensuring high-quality AIGC services in the metaverse setting.
IV-C2 Framework configuration
In our experiments, we call the GPT-4 model through the OpenAI API to implement the pluggable LLM module, and the RAG module is built on top of LangChain [^6] https://www.langchain.com/. We set the chunk size, chunk overlap, and retrieval results are set as $10001000$, $200200$, and $44$, respectively. Thus, the LLM agent can generate accurate models with a minimum number of retrieved tokens, i.e., a total of $40004000$.